### v1

# A Simple Algorithm for Trimmed Multipoint Evaluation

Abstract
Evaluating a polynomial on a set of points is a fundamental task in computer algebra. In this work, we revisit a particular variant called trimmed multipoint evaluation: given an $n$-variate polynomial with bounded individual degree $d$ and total degree $D$, the goal is to evaluate it on a natural class of input points. This problem arises as a key subroutine in recent algorithmic results [Dinur; SODA ’21], [Dell, Haak, Kallmayer, Wennmann; SODA ’25]. It is known that trimmed multipoint evaluation can be solved in near-linear time [van der Hoeven, Schost; AAECC ’13] by a clever yet somewhat involved algorithm. We give a simple recursive algorithm that avoids heavy computer-algebraic machinery, and can be readily understood by researchers without specialized background.
1 Introduction
One of the most fundamental problems in computer algebra is to efficiently evaluate a polynomial $P$ on some set of points, known as the multipoint evaluation problem. Besides its importance as one of the most basic algebraic primitives, this problem finds many further applications in computer algebra (such as modular composition and polynomial factorization) and in algorithm design beyond (in diverse fields such as computational geometry, coding theory and cryptography). The inverse task, to interpolate a polynomial from a given set of evaluations, is an equally important primitive.
For univariate polynomials, a textbook algorithm [10, 26] solves the multipoint evaluation problem in near-linear time. This algorithm generalizes to multivariate polynomials [20] (via a simple divide-and-conquer method sometimes referred to as Yates’ algorithm [28]), however, only in the restricted setting where the evaluation points form a cartesian grid. Specifically, Yates’ algorithm evaluates an $n$-variate polynomials with individual degree $d$ on all points of a grid
$Z=\{z_{1,0},\dots,z_{1,d}\}\times\dots\times\{z_{n,0},\dots,z_{n,d}\}$
in near-optimal time [^1] Here and throughout, we write $O^{*}(\cdot)$ to omit polynomial factors in $n$ and $d$. $O^{*}((d+1)^{n})$. Lifting this restriction to grid points has been the focus of a long and active line of research [18, 23, 15, 25, 6, 24, 2, 1] which, following breakthroughs by Umans [23] and Kedlaya and Umans [15], only recently culminated in an algorithm with almost-optimal running time $(d+1)^{(1+o(1))n}\operatorname{poly}(n,d,\log|\mathbb{F}|)$, for all finite fields $\mathbb{F}$ and for $(d+1)^{n}$ arbitrary evaluation points, due to Bhargava, Ghosh, Guo, Kumar and Umans [1]. This fully settles the multipoint evaluation problem for dense polynomials (over finite fields), but leaves open whether almost-linear time can also be achieved for (some classes of) sparse polynomials and evaluation points.
In this paper we focus on one natural such class, called trimmed multipoint evaluation, with important applications in the design of exact and parameterized algorithms. Trimmed multipoint evaluation can be solved in near-linear time by an algorithm due to van der Hoeven and Schost [25]. Our contribution is that we make this result accessible to modern algorithm design (beyond computer algebra) by distilling a particularly simple recursive algorithm.
Trimmed Multipoint Evaluation
In the trimmed multipoint evaluation problem we focus on the class of $n$-variate polynomials $P$ with individual degree $d$ and total degree [^2] Recall that the individual degree $d$ of a polynomial is the largest exponent of a variable in a monomial, whereas the total degree $D$ is the largest sum of exponents in a monomial. E.g., $X_{1}^{2}X_{2}$ has individual degree $d=2$ and total degree $D=3$. $D$. This is a very natural class of polynomials which, for $D\ll nd$, is exponentially sparser than polynomials with just an individual degree bound. As evaluation points we consider triangular subsets of grids $Z$ defined by
$\{(z_{1,\ell_{1}},\dots,z_{n,\ell_{n}}):\ell\in\{0,\dots,d\}^{n},\,\ell_{1}+\dots+\ell_{n}\leq D\}\subseteq Z,$
to which we will informally refer as trimmed grids. This is arguably the most naturally matching class of evaluation points. To see this, first observe that the number of relevant (i.e., possibly nonzero) coefficients of $P$ equals exactly the number of grid points. We denote this number by $\binom{n}{\leq D}_{d}$ (which can be seen as an appropriate generalization of a binomial coefficient called an extended binomial coefficient). More importantly, it turns out that the polynomial $P$ is uniquely determined by the evaluations on an appropriate trimmed grid, i.e., we can interpolate $P$ given only these evaluations. Trimmed grids are, in a sense, the only sets of evaluation points satisfying this property; see [22].
Applications
Besides being a natural problem in its own right, our interest in trimmed multipoint evaluation stems mainly from its applications in the context of exponential-time algorithms, most notably in a line of research on solving systems of polynomial equations [17, 7, 13, 12]. In this problem the input consists of $n$-variate polynomials $P_{1},\dots,P_{m}$ over some finite field $\mathbb{F}_{q}$ with (total) degree $\Delta$, and the task is to test if all polynomials simultaneously vanish at some point $x\in\mathbb{F}_{q}^{n}$, i.e., $P_{1}(x)=\dots=P_{m}(x)=0$. The initial breakthrough due to Lokshtanov, Paturi, Tamaki, Williams and Yu [17] established that this problem can be solved exponentially faster than brute-force, in time $(q-\epsilon)^{n}$ for some $\epsilon>0$, whenever $q$ and $\Delta$ are constant. This inspired several follow-up papers aiming to optimize the precise exponential running time [7, 13, 12]. Trimmed multipoint evaluation shows up as a critical subroutine in the algorithms due to Dinur [13] (for multilinear polynomials, $d=q-1=1$) and due to Dell, Haak, Kallmayer and Wennmann [12] (for general $d=q-1\geq 1$). Solving systems of polynomial equations in turn has many more applications on both the theoretical side—e.g., parity-counting directed Hamiltonian cycles [4]—and the practical side—e.g., the security of several so-called multivariate cryptosystems is based on the hardness of solving quadratic equations [13, 21, 16]. Especially for the latter it could be interesting to achieve simple, implementable algorithms.
In another closely related work, Björklund, Husfeldt, Kaski and Koivisto [5] considered trimmed [^3] In fact, they consider a more general definition of “trimmed” allowing arbitrary downward-closed sets. variants of the Zeta and Möbius transforms to develop fast algorithms for various exponential-time graph problems, such as computing the chromatic number for constant-degree-bounded graphs. The Zeta and Möbius transforms can be regarded as special cases of polynomial multipoint evaluation. [^4] Indeed, recall that the zeta transform of a function $f:\{0,1\}^{n}\to\mathbb{F}$ is defined as $(f\zeta)(X)=\sum_{Y\subseteq X}f(Y)$, where we identify sets $X\subseteq[n]$ with their indicator vectors $X\in\{0,1\}^{n}$. Consider the $n$-variate polynomial $P(x_{1},\dots,x_{n})=\sum_{Y\subseteq[n]}f(Y)\prod_{i\in Y}x_{i}$, and observe that $(f\zeta)(X)=P(X)$. Thus, the zeta transform $f\zeta$ can be read off the evaluations of $P$ on the grid $\{0,1\}^{n}$.
State of the Art
The state of the art for trimmed multipoint evaluation, as mentioned before, is a clever algorithm with near-linear running time $O^{*}(\binom{n}{\leq D}_{d})$ due to van der Hoeven and Schost [25]. Their algorithm is based on the classical concept of Newton interpolation (see e.g. [3]), suitably tailored to the trimmed problem (see also [27, 9, 22, 11, 14] for some more references with a more mathematical point of view). The algorithm also offers two additional benefits: (1) It even solves a strictly more general multipoint evaluation problem on arbitrary downward-closed sets of relevant coefficients and grid points. (2) van der Hoeven and Schost have optimized the lower-order factors achieving an algebraic algorithm with $O(nN\log^{2}N\log\log N)$ field operations where $N=\binom{n}{\leq D}_{d}$. The potentially remaining lower-order improvements even persist in the univariate setting. Thus, all in all, the trimmed multipoint evaluation problem has already been satisfyingly resolved.
The only downside is that van der Hoeven and Schost’s algorithm is arguably somewhat intricate—both in the sense that it can be technically demanding to understand, particularly for researchers outside the computer algebra community, and in that it relies on several textbook algebraic primitives, such as efficient conversions between polynomial bases [3] and the use of truncated Fourier transforms as an implementation detail to achieve further improvements.
Our Contribution
Our focus here is not to optimize the running time or generality of this state of the art. Instead, given the many exciting algorithmic applications our contribution is to make van der Hoeven and Schost’s result for trimmed multipoint evaluation accessible to the algorithms community. We design a simple recursive algorithm that is teachable to researchers without any background in computer algebra. Moreover, our algorithm does not rely on any black-box algebraic primitives other than Gaussian elimination. As in [25], we also obtain an equally simple algorithm for the interpolation problem.
Our emphasis on simplicity clashes, however, with the additional benefits (1) and (2): (1) It seems hard to obtain a recursive algorithm for the more general problem, and (2) optimizing lower-order factors would involve dealing with more details. Besides, for exponential-time algorithms we are typically anyway not bothered with keeping track of polynomial factors. For these reasons we have decided to stick to the simplest version, resulting in a pleasingly simple 8-line algorithm.
Remark on Terminology
The term “trimmed” multipoint evaluation is inspired by the related algorithms for Zeta and Möbius transforms in [5]. It is not standard in the mathematical literature, where one more often encounters terms like “triangular subsets of tensor product grids”. We adopt the “trimmed” terminology here as this paper is primarily intended for the algorithms community.
2 Preliminaries
We write $[n]=\{1,\dots,n\}$. Throughout, let $\mathbb{F}$ be a field and assume that we can evaluate field operations in unit time. For integers $n,k,d$ with $n\geq 0$ and $d\geq 1$ we define the extended binomial coefficient
$\binom{n}{k}_{d}=\lvert\{\ell\in\{0,\dots,d\}^{n}:\ell_{1}+\dots+\ell_{n}=k\}\rvert.$
That is, $\binom{n}{k}_{d}$ counts the number of multisubsets of $[n]$ with size $k$ and multiplicity at most $d$, or equivalently, the number of monomials with total degree $k$ and individual degree at most $d$ in an $n$-variate polynomial. In the same spirit we define the set
$\binom{[n]}{k}_{d}=\{\ell\in\{0,\dots,d\}^{n}:\ell_{1}+\dots+\ell_{n}=k\}.$
As a shorthand, we write $\binom{n}{\leq k}_{d}=\sum_{i=0}^{k}\binom{n}{i}_{d}$, and similarly define $\binom{[n]}{\leq k}_{d}=\bigcup_{i=0}^{k}\binom{n}{i}_{d}$. Further, we rely on the following generalization of Pascal’s triangle; see [8].
Lemma 2.1 (Extended Pascal Triangle).
For $n,d\geq 1$ it holds that
$\binom{n}{k}_{d}=\sum_{j=0}^{d}\binom{n-1}{k-j}_{d}.$
Throughout we refer to a set $Z=\{z_{1,0},\dots,z_{1,d}\}\times\dots\times\{z_{n,0},\dots,z_{n,d}\}$ of field elements $z_{i,j}$ as a grid. To conveniently refer to the grid points we regularly write $Z_{\ell}=(z_{1,\ell_{1}},\dots,z_{n,\ell_{n}})$ for $\ell\in\{0,\dots,d\}^{n}$. In particular, the subset of grid points $Z_{\ell}$ where $\ell$ ranges over $\binom{[n]}{\leq D}_{d}$ constitutes exactly a trimmed grid as introduced before.
3 Trimmed Multipoint Evaluation and Interpolation
In this section, we present a simple, algebraic algorithm for trimmed multipoint evaluation. As it is usually the case, the same algorithmic approach yields a simple algorithm for trimmed interpolation as well.
3.1 Key Ideas
The problem of evaluating univariate polynomials can be defined via the Vandermonde matrix. For $z_{0},\dots,z_{d}\in\mathbb{F}$, define the (square) Vandermonde matrix $V\in\mathbb{F}^{(d+1)\times(d+1)}$ as
$V=V(z_{0},\dotsc,z_{d})=\begin{pmatrix}1&z_{0}&z_{0}^{2}&\cdots&z_{0}^{d}\\ 1&z_{1}&z_{1}^{2}&\cdots&z_{1}^{d}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&z_{d}&z_{d}^{2}&\cdots&z_{d}^{d}\end{pmatrix}.$
Let $a\in\mathbb{F}^{d+1}$ be the coefficient vector of the univariate polynomial $P(X)=\sum_{i=0}^{d}a_{i}X^{i}$ of degree $d$. The evaluation of $P$ at all points $z_{j}$ with $j\in\{0,\dotsc,d\}$ can be expressed as the matrix-vector product $V\cdot a=y$ (indeed, each entry is of the form $y_{j}=\sum_{i=0}^{d}a_{i}z_{j}^{i}=P(z_{j})$). Similarly, we can interpolate $P$ from its evaluations $P(z_{j})$ using the matrix-vector product $V^{-1}\cdot y=V^{-1}\cdot V\cdot a=a$. The generalization to multivariate polynomials is simple: Let $a\in\mathbb{F}^{(d+1)^{n}}$ be the coefficient vector of an $n$-variate polynomial $P$ with individual degree $d$ that we want to evaluate on all grid points in $Z$. Then the Kronecker product $V(z_{1,0},\dotsc,z_{1,d})\otimes\cdots\otimes V(z_{n,0},\dotsc,z_{n,d})\cdot a=y$ yields the vector $y\in\mathbb{F}^{(d+1)^{n}}$ of all grid point evaluations. The classical algorithm of Yates [28] allows to compute these evaluations recursively. Notably, this computation does not account for the total degree $D$ and automatically results in a running time of $(d+1)^{n}$ for both multipoint evaluation and interpolation—even in our setting where the total degree $D$ is much smaller than its maximal value $nd$. Consequently, we need more insights to tailor this approach to the trimmed requirements.
Idea 1: Recursion Scheme.
Our goal is to achieve a running time that is linear in $\binom{n}{\leq D}_{d}$, i.e., in the number of grid points on which we want to evaluate a multivariate polynomial $P$. As a baseline, we start with a short explanation of Yates’ algorithm: write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{i}$ of degree $D-i$, where each $P_{i}$ can be seen as a coefficient of a univariate polynomial in $X_{n}$. By making $d+1$ many recursive calls of size $\binom{n-1}{\leq D}_{d}$, we compute the evaluations $P_{i}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$. We obtain all $P(Z_{\ell})$ for $\ell\in\binom{[n]}{\leq D}_{d}$ by evaluating the univariate polynomials $\sum_{i=0}^{d}P_{i}(Z_{\ell^{\prime}})\cdot X_{n}^{i}$ at all grid points $z_{n,i}$. However, this does not exploit the degrees $D-i$ of the polynomials $P_{i}$ in the recursive calls, and hence results in the running time of $(d+1)^{n}$.
Consider the following identity of the extended binomial coefficient that can be derived from Lemma 2.1
$\binom{n}{\leq D}_{d}=\binom{n-1}{\leq D}_{d}+\binom{n-1}{\leq D-1}_{d}+\cdots+\binom{n-1}{\leq D-d}_{d}.$
In light of this identity, in order to achieve a running time of $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$ we aim to design an algorithm with $d+1$ recursive calls on $n-1$ variables with total degrees $D,D-1,\dotsc,D-d$.
1
Input : $n$-variate polynomial $P$ of individual degree $d$ and total degree $D$
Output : Evaluations $P(Z_{\ell})$ for all $\ell\in\binom{[n]}{\leq D}_{d}$
2
3
4if $n=0$ then return the constant $P$.
5
6Write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$.
7Compute the LU decomposition $V(z_{n,0},\dots,z_{n,d})=L\cdot U$.
8Compute the vector of polynomials
$\hskip 85.35826pt\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,d}\\ &\ddots&\vdots\\ &&U_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}.$
9for $j=0,\dots,d$  do
10       Recursively call TrimmedEvaluation$(Q_{j})$ to evaluate $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
11
12for $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$ do
13       Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
$\hskip 34.14322pt\begin{pmatrix}P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen\\ \vdots\\ P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{k,0}&\cdots&L_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}(Z_{\ell^{\prime}})\\ \vdots\\ Q_{k}(Z_{\ell^{\prime}})\\ \end{pmatrix}.$ ) .
Algorithm 1 TrimmedEvaluation$(P)$
Idea 2: LU Decomposition.
Our goal is an interleaving of the $P_{i}$’s into polynomials $Q_{j}$ satisfying two properties: (1) All evaluations $Q_{j}(Z_{\ell^{\prime}})$ can be recursively computed in a call of size $\binom{n-1}{\leq D-j}_{d}$; in particular, $Q_{j}$ must have degree $D-j$. (2) Simultaneously, we need to recover all evaluations $P(Z_{\ell})$ from the recursively computed evaluations $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$. Due to the degree restriction for the polynomials $Q_{j}$ in (1) and evaluations $Q_{j}(Z_{\ell^{\prime}})$ in (2), this coincides with a matrix factorization of $V=L\cdot U$ where $U$ is upper-triangular and $L$ lower-triangular, i.e., a LU decomposition.
In more detail: Why is the LU decomposition useful in keeping the degrees of the polynomials $Q_{j}$ “small”? Let $p=(P_{0},\dots,P_{d})^{T}$ be the vector of polynomials $P_{i}$ and consider the product $U\cdot p=(Q_{1},\dotsc,Q_{d})^{T}=q$. Here, the upper triangular structure of $U$ guarantees that each $Q_{j}=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}$ has at most degree $D-j$, since the first non-zero entry in the $j$-th row of $U$ corresponds to polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree. Additionally, we can also recover $P(Z_{\ell})$ from only the recursively computed $Q_{j}(Z_{\ell^{\prime}})$ as $L$ has lower triangular structure—appropriately exploiting the fact that $L\cdot q=L\cdot U\cdot p=V\cdot p$ [^5] The analogy to van der Hoeven and Schost’s algorithm is as follows: Broadly speaking, their algorithm is based on efficient transformations between not only the coefficient and evaluation-based representations of polynomials, but also involving a third representation based on so-called Newton polynomials. In this terminology, the matrix $U$ performs a basis change from the monomial basis to the Newton basis, and the matrix $L$ performs a basic change from the Newton basis to the evaluation basis. With these ideas in mind, we are now in the position to present the algorithm in detail.
3.2 Trimmed Multipoint Evaluation
Throughout this section, we will prove the following theorem.
Theorem 3.1 (Trimmed Multipoint Evaluation).
Let $P$ be an $n$-variate polynomial over $\mathbb{F}$ with individual degree $d$ and total degree $D$, and let $Z$ be a grid. The evaluations $P(Z_{\ell})$ can be computed for all $\ell\in\binom{[n]}{\leq D}_{d}$ in time $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$.
For the proof of Theorem 3.1, consider the algorithm TrimmedEvaluation. Throughout we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
Lemma 3.2 (Correctness of TrimmedEvaluation).
Given an $n$-variate polynomial $P$ over $\mathbb{F}$ with individual degree $d$ and total degree $D$, TrimmedEvaluation correctly computes the evaluations $P(Z_{\ell})$ for all $\ell\in\binom{[n]}{\leq D}_{d}$.
Proof 3.3.
Given a polynomial $P$ and a grid $Z$, we show that TrimmedEvaluation correctly computes all evaluations $P(Z_{\ell})$ by induction on $n$.
For $n=0$, the polynomial $P$ does not depend on any variable. Thus, we correctly return the constant $P$. For $n\geq 1$, we write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{0},\dots,P_{d}$ of degrees $D,\dots,D-d$, respectively. Formally, each polynomial $P_{i}$ is the coefficient of $P$ viewed as a univariate polynomial in $X_{n}$. Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$ be a Vandermonde matrix. It is well-known that $V$ is invertible if and only if all grid points $z_{n,i}$ are distinct, and in this case the LU decomposition $V=L\cdot U$ in Line 1 always exists (see e.g. [19]). Further, let $Q_{0},\dotsc,Q_{d}$ as computed in Line 8 of TrimmedEvaluation. Note that each polynomial $Q_{j}(X_{1},\dotsc,X_{n-1})=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}(X_{1},\dotsc,X_{n-1})$ has degree $D-j$. Indeed, the upper triangular structure of $U$ guarantees that $Q_{j}$ has at most degree $D-j$, because the first non-zero entry in the $j$-th row of $U$ corresponds to the polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
Calling TrimmedEvaluation$(Q_{j})$, we recursively compute the evaluations $Q_{j}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ in Lines 1 and 1. Next, we focus on any iteration $\ell^{\prime}$ of the loop in Line 1. Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$, then we show that we correctly compute the evaluations $P(Z_{\ell})$. For each $j\in\{0,\dotsc,k\}$, we have
$\sum_{i=0}^{k}L_{j,i}\cdot Q_{i}(Z_{\ell^{\prime}})$ $=\sum_{i=0}^{d}L_{j,i}\cdot Q_{i}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}L_{j,i}\cdot\sum_{m=0}^{d}U_{i,m}\cdot P_{m}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}z_{n,j}^{i}\cdot P_{i}(Z_{\ell^{\prime}})$
$=P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen,$
where the first equality follows from the fact that $L_{j,i}=0$ whenever $i>k\geq j$. Notably, restricting the computation to $k$ is crucial to ensure that we only use the evaluations $Q_{j}(Z_{\ell^{\prime}})$ that we actually computed. This proves that TrimmedEvaluation correctly computes the evaluations $P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen,\dotsc,P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen$ ) in Line 13. Since each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ as above, the loop in Line 1 recovers all evaluations $P(Z_{\ell})$. As a result, the algorithm TrimmedEvaluation correctly computes all evaluations $P(Z_{\ell})$.
Lastly, we prove the running time of TrimmedEvaluation.
Lemma 3.4 (Running Time of TrimmedEvaluation).
The algorithm TrimmedEvaluation runs in time $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$.
Proof 3.5.
Let $T(n,d,D)$ be the running time of TrimmedEvaluation. The time to compute the LU decomposition in Line $4$ and the matrix-vector products in Lines $5$ and $9$ can be bounded by $\operatorname{poly}(n,d)$. Consequently, the running time of the algorithm without the recursive calls in Line $7$ is $\binom{n}{\leq D}_{d}\cdot M(n,d)$ for some function [^6] In our context, we are content with bounding $M(n,d)=\operatorname{poly}(n,d)$; however, let us remark that following van der Hoeven and Schost [25] we could achieve a dependence on $d$ that is only polylogarithmic. The main insight is that the L- and U-factors of a Vandermonde matrix are sufficiently structured to support matrix-vector operations in time $O(d(\log d)^{O(1)})$. $M(n,d)=\operatorname{poly}(n,d)$. Thus, the algorithm admits the following recurrence
$T(n,d,D)$ $\leq\binom{n}{\leq D}_{d}\cdot M(n,d)+\sum_{i=0}^{d}T(n-1,d,D-i).$
By induction on $n$, we show that $T(n,d,D)\leq\binom{n}{\leq D}_{d}\cdot M(n,d)\cdot n$. Indeed, it holds that
$T(n,d,D)$ $\leq\binom{n}{\leq D}_{d}\cdot M(n,d)+\sum_{i=0}^{d}T(n-1,d,D-i)$
$\leq\binom{n}{\leq D}_{d}\cdot M(n,d)+\sum_{i=0}^{d}\binom{n}{\leq D-i}_{d}\cdot M(n,d)\cdot(n-1)$
$=\binom{n}{\leq D}_{d}\cdot M(n,d)+\sum_{i=0}^{d}\sum^{D-i}_{j=0}\binom{n-1}{j}_{d}\cdot M(n,d)\cdot(n-1)$
$\leq\binom{n}{\leq D}_{d}\cdot M(n,d)+\sum^{D}_{j=0}\binom{n}{j}_{d}\cdot M(n,d)\cdot(n-1)$
$=\binom{n}{\leq D}_{d}\cdot M(n,d)+\binom{n}{\leq D}_{d}\cdot M(n,d)\cdot(n-1)$
$=\binom{n}{\leq D}_{d}\cdot M(n,d)\cdot n$
Therefore, TrimmedEvaluation runs in time $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$.
Combining Lemma 3.2 and Lemma 3.4 concludes the proof of Theorem 3.1.
1
Input : Evaluations $\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$
Output : Unique $n$-variate polynomial $P$ of individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$
2
3
4if $n=0$ then return the constant $P$.
5 Compute the LU decomposition $V(z_{n,0},\dots,z_{n,d})^{-1}=L\cdot U$.
6 for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ do
7       Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
$\hskip 35.56593pt\begin{pmatrix}\beta_{(\ell^{\prime},0)}\\ \vdots\\ \beta_{(\ell^{\prime},k)}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,k}\\ &\ddots&\vdots\\ &&U_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}\alpha_{(\ell^{\prime},0)}\\ \vdots\\ \alpha_{(\ell^{\prime},k)}\\ \end{pmatrix}.\;$
8
9for $j=0,\dots,d$ do
10       Recursively call TrimmedInterpolation($\left\lparen(\beta_{(\ell^{\prime},j)})_{\ell^{\prime}}\right\rparen$ ) for ${\ell^{\prime}}\in\binom{[n-1]}{\leq D-j}_{d}$ to interpolate the $(n-1)$-variate polynomial $Q_{j}$.
11
12
13Compute the vector of polynomials
$\hskip 65.44142pt\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{d,0}&\cdots&L_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}.\;$
14Compute $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$.
Algorithm 2 TrimmedInterpolation$((\alpha_{\ell})_{\ell})$
3.3 Trimmed Interpolation
Throughout this section, we show that a polynomial can be (uniquely) interpolated from its evaluations on the trimmed grid points—using essentially the same algorithmic approach as in Section 3.2.
Theorem 3.6 (Trimmed Interpolation).
Let $\alpha_{\ell}\in\mathbb{F}$ for $\ell\in\binom{[n]}{\leq D}_{d}$, and let $Z$ be a grid. The unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ that satisfies $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$ can be computed in time $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$.
Consider the algorithm TrimmedInterpolation for the proof of Theorem 3.6. As before, we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
Lemma 3.7 (Correctness of TrimmedInterpolation).
Given evaluations $\alpha_{\ell}$ for $\ell\in\binom{[n]}{\leq D}_{d}$, TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$.
Proof 3.8.
Given the evaluations $\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$, we show that TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ by induction on $n$. If $n=0$, simply return the constant $P$ as it does not depend on any variable.
For the remainder of the proof we assume $n\geq 1$. Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$. It is well-known that $V$ is invertible if and only if the grid points $z_{n,0},\dots,z_{n,d}$ are distinct. Additionally, it implies that the LU decomposition in Line 2, namely $V^{-1}=L\cdot U$, always exists (see e.g. [19]).
In the following, we show that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$. Note that each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ with $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ and $\sum_{i=1}^{n-1}\ell^{\prime}_{i}+j\leq D$. Fix such a pair $\ell^{\prime}$ and $j$, and let $k=\min\{d,D-\sum_{i=1}^{n-1}\ell^{\prime}_{i}\}$. It holds that
$P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen$ ) $=\sum_{i=0}^{d}z_{n,j}^{i}\cdot P_{i}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\beta_{(\ell^{\prime},m)}$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\sum_{h=0}^{k}U_{m,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{h=0}^{k}(V^{-1})_{i,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\sum_{h=0}^{k}I_{j,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\alpha_{(\ell^{\prime},j)},$
where $I\in\mathbb{F}^{{d+1}\times{d+1}}$ is the identity matrix. The last equality uses the fact that $j\leq k$. Consequently, TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$.
As the recursion scheme of TrimmedInterpolation and TrimmedEvaluation are exactly the same, we refer to Lemma 3.4 for a proof of the running time of TrimmedInterpolation. This concludes the proof of Theorem 3.6.

### v2

@@
 We give a simple recursive algorithm that avoids heavy computer-algebraic machinery, and can be readily understood by researchers without specialized background.
+keywords:
+Algebraic Algorithms, Multipoint Evaluation, Interpolation, LU Decomposition
 1 Introduction
@@
 The inverse task, to interpolate a polynomial from a given set of evaluations, is an equally important primitive.
-For univariate polynomials, a textbook algorithm [10, 26] solves the multipoint evaluation problem in near-linear time.
-This algorithm generalizes to multivariate polynomials [20] (via a simple divide-and-conquer method sometimes referred to as Yates’ algorithm [28]), however, only in the restricted setting where the evaluation points form a cartesian grid.
+For univariate polynomials, a textbook algorithm [BorodinM74, vonzurGathenG13] solves the multipoint evaluation problem in near-linear time.
+This algorithm generalizes to multivariate polynomials [Pan94] (via a simple divide-and-conquer method sometimes referred to as Yates’ algorithm [Yates37]), however, only in the restricted setting where the evaluation points form a cartesian grid.
 Specifically, Yates’ algorithm evaluates an $n$-variate polynomials with individual degree $d$ on all points of a grid
@@
 $O^{*}((d+1)^{n})$.
-Lifting this restriction to grid points has been the focus of a long and active line of research [18, 23, 15, 25, 6, 24, 2, 1] which, following breakthroughs by Umans [23] and Kedlaya and Umans [15], only recently culminated in an algorithm with almost-optimal running time $(d+1)^{(1+o(1))n}\operatorname{poly}(n,d,\log|\mathbb{F}|)$, for all finite fields $\mathbb{F}$ and for $(d+1)^{n}$ arbitrary evaluation points, due to Bhargava, Ghosh, Guo, Kumar and Umans [1].
+Lifting this restriction to grid points has been the focus of a long and active line of research [NuskenZ04, Umans08, KedlayaU11, HoevenS13, BjorklundKW19, HoevenL21a, BhargavaG0M22, BhargavaGG0U22] which, following breakthroughs by Umans [Umans08] and Kedlaya and Umans [KedlayaU11], only recently culminated in an algorithm with almost-optimal running time $(d+1)^{(1+o(1))n}\operatorname{poly}(n,d,\log|\mathbb{F}|)$, for all finite fields $\mathbb{F}$ and for $(d+1)^{n}$ arbitrary evaluation points, due to Bhargava, Ghosh, Guo, Kumar and Umans [BhargavaGG0U22].
 This fully settles the multipoint evaluation problem for dense polynomials (over finite fields), but leaves open whether almost-linear time can also be achieved for (some classes of) sparse polynomials and evaluation points.
 In this paper we focus on one natural such class, called trimmed multipoint evaluation, with important applications in the design of exact and parameterized algorithms.
-Trimmed multipoint evaluation can be solved in near-linear time by an algorithm due to van der Hoeven and Schost [25].
+Trimmed multipoint evaluation can be solved in near-linear time by an algorithm due to van der Hoeven and Schost [HoevenS13].
 Our contribution is that we make this result accessible to modern algorithm design (beyond computer algebra) by distilling a particularly simple recursive algorithm.
@@
 More importantly, it turns out that the polynomial $P$ is uniquely determined by the evaluations on an appropriate trimmed grid, i.e., we can interpolate $P$ given only these evaluations.
-Trimmed grids are, in a sense, the only sets of evaluation points satisfying this property; see [22].
+Trimmed grids are, in a sense, the only sets of evaluation points satisfying this property; see [Sauer04].
 Applications
-Besides being a natural problem in its own right, our interest in trimmed multipoint evaluation stems mainly from its applications in the context of exponential-time algorithms, most notably in a line of research on solving systems of polynomial equations [17, 7, 13, 12].
+Besides being a natural problem in its own right, our interest in trimmed multipoint evaluation stems mainly from its applications in the context of exponential-time algorithms, most notably in a line of research on solving systems of polynomial equations [LokshtanovPTWY17, BjorklundK019, Dinur21, DellHKW25].
 In this problem the input consists of $n$-variate polynomials $P_{1},\dots,P_{m}$ over some finite field $\mathbb{F}_{q}$ with (total) degree $\Delta$, and the task is to test if all polynomials simultaneously vanish at some point $x\in\mathbb{F}_{q}^{n}$, i.e., $P_{1}(x)=\dots=P_{m}(x)=0$.
-The initial breakthrough due to Lokshtanov, Paturi, Tamaki, Williams and Yu [17] established that this problem can be solved exponentially faster than brute-force, in time $(q-\epsilon)^{n}$ for some $\epsilon>0$, whenever $q$ and $\Delta$ are constant.
-This inspired several follow-up papers aiming to optimize the precise exponential running time [7, 13, 12].
-Trimmed multipoint evaluation shows up as a critical subroutine in the algorithms due to Dinur [13] (for multilinear polynomials, $d=q-1=1$) and due to Dell, Haak, Kallmayer and Wennmann [12] (for general $d=q-1\geq 1$).
-Solving systems of polynomial equations in turn has many more applications on both the theoretical side—e.g., parity-counting directed Hamiltonian cycles [4]—and the practical side—e.g., the security of several so-called multivariate cryptosystems is based on the hardness of solving quadratic equations [13, 21, 16].
+The initial breakthrough due to Lokshtanov, Paturi, Tamaki, Williams and Yu [LokshtanovPTWY17] established that this problem can be solved exponentially faster than brute-force, in time $(q-\epsilon)^{n}$ for some $\epsilon>0$, whenever $q$ and $\Delta$ are constant.
+This inspired several follow-up papers aiming to optimize the precise exponential running time [BjorklundK019, Dinur21, DellHKW25].
+Trimmed multipoint evaluation shows up as a critical subroutine in the algorithms due to Dinur [Dinur21] (for multilinear polynomials, d=q1=1d=q-1=1) and due to Dell, Haak, Kallmayer and Wennmann [DellHKW25] (for general $d=q-1\geq 1$).
+Solving systems of polynomial equations in turn has many more applications on both the theoretical side—e.g., parity-counting directed Hamiltonian cycles [BjorklundH13]—and the practical side—e.g., the security of several so-called multivariate cryptosystems is based on the hardness of solving quadratic equations [Dinur21, Patarin96, KipnisPG99].
 Especially for the latter it could be interesting to achieve simple, implementable algorithms.
-In another closely related work, Björklund, Husfeldt, Kaski and Koivisto [5] considered trimmed [^3] In fact, they consider a more general definition of “trimmed” allowing arbitrary downward-closed sets.
+In another closely related work, Björklund, Husfeldt, Kaski and Koivisto [BjorklundHKK10] considered trimmed [^3] In fact, they consider a more general definition of “trimmed” allowing arbitrary downward-closed sets.
 variants of the Zeta and Möbius transforms to develop fast algorithms for various exponential-time graph problems, such as computing the chromatic number for constant-degree-bounded graphs.
@@
 State of the Art
-The state of the art for trimmed multipoint evaluation, as mentioned before, is a clever algorithm with near-linear running time $O^{*}(\binom{n}{\leq D}_{d})$ due to van der Hoeven and Schost [25].
+The state of the art for trimmed multipoint evaluation, as mentioned before, is a clever algorithm with near-linear running time $O^{*}(\binom{n}{\leq D}_{d})$ due to van der Hoeven and Schost [HoevenS13].
 Their algorithm is based on the classical concept of Newton interpolation (see e.g.
-[3]), suitably tailored to the trimmed problem (see also [27, 9, 22, 11, 14] for some more references with a more mathematical point of view).
+[BiniP94]), suitably tailored to the trimmed problem (see also [Werner80, Boor92, Sauer04, ChkifaCS14, DynF14] for some more references with a more mathematical point of view).
 The algorithm also offers two additional benefits: (1) It even solves a strictly more general multipoint evaluation problem on arbitrary downward-closed sets of relevant coefficients and grid points.
@@
 Thus, all in all, the trimmed multipoint evaluation problem has already been satisfyingly resolved.
-The only downside is that van der Hoeven and Schost’s algorithm is arguably somewhat intricate—both in the sense that it can be technically demanding to understand, particularly for researchers outside the computer algebra community, and in that it relies on several textbook algebraic primitives, such as efficient conversions between polynomial bases [3] and the use of truncated Fourier transforms as an implementation detail to achieve further improvements.
+The only downside is that van der Hoeven and Schost’s algorithm is arguably somewhat intricate—both in the sense that it can be technically demanding to understand, particularly for researchers outside the computer algebra community, and in that it relies on several textbook algebraic primitives, such as efficient conversions between polynomial bases [BiniP94] and the use of truncated Fourier transforms as an implementation detail to achieve further improvements.
 Our Contribution
@@
 Moreover, our algorithm does not rely on any black-box algebraic primitives other than Gaussian elimination.
-As in [25], we also obtain an equally simple algorithm for the interpolation problem.
+As in [HoevenS13], we also obtain an equally simple algorithm for the interpolation problem.
 Our emphasis on simplicity clashes, however, with the additional benefits (1) and (2): (1) It seems hard to obtain a recursive algorithm for the more general problem, and (2) optimizing lower-order factors would involve dealing with more details.
@@
 Remark on Terminology
-The term “trimmed” multipoint evaluation is inspired by the related algorithms for Zeta and Möbius transforms in [5].
+The term “trimmed” multipoint evaluation is inspired by the related algorithms for Zeta and Möbius transforms in [BjorklundHKK10].
 It is not standard in the mathematical literature, where one more often encounters terms like “triangular subsets of tensor product grids”.
@@
 2 Preliminaries
-We write $[n]=\{1,\dots,n\}$.
+We write $n]=\{1,\dots,n\}$.
 Throughout, let $\mathbb{F}$ be a field and assume that we can evaluate field operations in unit time.
@@
 As a shorthand, we write $\binom{n}{\leq k}_{d}=\sum_{i=0}^{k}\binom{n}{i}_{d}$, and similarly define $\binom{[n]}{\leq k}_{d}=\bigcup_{i=0}^{k}\binom{n}{i}_{d}$.
-Further, we rely on the following generalization of Pascal’s triangle; see [8].
+Further, we rely on the following generalization of Pascal’s triangle; see [Bollinger93].
 Lemma 2.1 (Extended Pascal Triangle).
@@
 Then the Kronecker product $V(z_{1,0},\dotsc,z_{1,d})\otimes\cdots\otimes V(z_{n,0},\dotsc,z_{n,d})\cdot a=y$ yields the vector $y\in\mathbb{F}^{(d+1)^{n}}$ of all grid point evaluations.
-The classical algorithm of Yates [28] allows to compute these evaluations recursively.
+The classical algorithm of Yates [Yates37] allows to compute these evaluations recursively.
 Notably, this computation does not account for the total degree $D$ and automatically results in a running time of $(d+1)^{n}$ for both multipoint evaluation and interpolation—even in our setting where the total degree $D$ is much smaller than its maximal value $nd$.
@@
 Our goal is to achieve a running time that is linear in $\binom{n}{\leq D}_{d}$, i.e., in the number of grid points on which we want to evaluate a multivariate polynomial $P$.
-As a baseline, we start with a short explanation of Yates’ algorithm: write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{i}$ of degree $D-i$, where each $P_{i}$ can be seen as a coefficient of a univariate polynomial in $X_{n}$.
+As a baseline, we start with a short explanation of Yates’ algorithm: write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many (n1)(n-1)-variate polynomials $P_{i}$ of degree DiD-i, where each $P_{i}$ can be seen as a coefficient of a univariate polynomial in $X_{n}$.
 By making $d+1$ many recursive calls of size $\binom{n-1}{\leq D}_{d}$, we compute the evaluations $P_{i}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$.
 We obtain all $P(Z_{\ell})$ for $\ell\in\binom{[n]}{\leq D}_{d}$ by evaluating the univariate polynomials $\sum_{i=0}^{d}P_{i}(Z_{\ell^{\prime}})\cdot X_{n}^{i}$ at all grid points $z_{n,i}$.
-However, this does not exploit the degrees $D-i$ of the polynomials $P_{i}$ in the recursive calls, and hence results in the running time of $(d+1)^{n}$.
-Consider the following identity of the extended binomial coefficient that can be derived from Lemma 2.1
+However, this does not exploit the degrees DiD-i of the polynomials $P_{i}$ in the recursive calls, and hence results in the running time of $(d+1)^{n}$.
+Consider the following identity of the extended binomial coefficient that can be derived from Lemma˜2.1
 $\binom{n}{\leq D}_{d}=\binom{n-1}{\leq D}_{d}+\binom{n-1}{\leq D-1}_{d}+\cdots+\binom{n-1}{\leq D-d}_{d}.$
-In light of this identity, in order to achieve a running time of $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$ we aim to design an algorithm with $d+1$ recursive calls on $n-1$ variables with total degrees $D,D-1,\dotsc,D-d$.
+In light of this identity, in order to achieve a running time of $O^{*}\big(\binom{n}{\leq D}_{d}\big)$ we aim to design an algorithm with $d+1$ recursive calls on n1n-1 variables with total degrees $D,D-1,\dotsc,D-d$.
 1
@@
 9for $j=0,\dots,d$ do
-10       Recursively call TrimmedEvaluation$(Q_{j})$ to evaluate $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
+[…]
 11
@@
 13   Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
-$\hskip 34.14322pt\begin{pmatrix}P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen\\ \vdots\\ P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{k,0}&\cdots&L_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}(Z_{\ell^{\prime}})\\ \vdots\\ Q_{k}(Z_{\ell^{\prime}})\\ \end{pmatrix}.$ ) .
-Algorithm 1 TrimmedEvaluation$(P)$
+[…]
+Algorithm 1 TrimmedEvaluation(P)(P)
 Idea 2: LU Decomposition.
-Our goal is an interleaving of the $P_{i}$’s into polynomials $Q_{j}$ satisfying two properties: (1) All evaluations $Q_{j}(Z_{\ell^{\prime}})$ can be recursively computed in a call of size $\binom{n-1}{\leq D-j}_{d}$; in particular, $Q_{j}$ must have degree $D-j$.
+Our goal is an interleaving of the $P_{i}$’s into polynomials $Q_{j}$ satisfying two properties: (1) All evaluations $Q_{j}(Z_{\ell^{\prime}})$ can be recursively computed in a call of size $\binom{n-1}{\leq D-j}_{d}$; in particular, $Q_{j}$ must have degree DjD-j.
 (2) Simultaneously, we need to recover all evaluations $P(Z_{\ell})$ from the recursively computed evaluations $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
@@
 Let $p=(P_{0},\dots,P_{d})^{T}$ be the vector of polynomials $P_{i}$ and consider the product $U\cdot p=(Q_{1},\dotsc,Q_{d})^{T}=q$.
-Here, the upper triangular structure of $U$ guarantees that each $Q_{j}=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}$ has at most degree $D-j$, since the first non-zero entry in the $j$-th row of $U$ corresponds to polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
-Additionally, we can also recover $P(Z_{\ell})$ from only the recursively computed $Q_{j}(Z_{\ell^{\prime}})$ as $L$ has lower triangular structure—appropriately exploiting the fact that $L\cdot q=L\cdot U\cdot p=V\cdot p$ [^5] The analogy to van der Hoeven and Schost’s algorithm is as follows: Broadly speaking, their algorithm is based on efficient transformations between not only the coefficient and evaluation-based representations of polynomials, but also involving a third representation based on so-called Newton polynomials.
+Here, the upper triangular structure of $U$ guarantees that each $Q_{j}=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}$ has at most degree DjD-j, since the first non-zero entry in the $j$-th row of $U$ corresponds to polynomial $P_{j}$ of degree DjD-j while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
+Additionally, we can also recover $P(Z_{\ell})$ from only the recursively computed $Q_{j}(Z_{\ell^{\prime}})$ as $L$ has lower triangular structure—appropriately exploiting the fact that $L\cdot q=L\cdot U\cdot p=V\cdot p$.
+[^5] The analogy to van der Hoeven and Schost’s algorithm is as follows: Broadly speaking, their algorithm is based on efficient transformations between not only the coefficient and evaluation-based representations of polynomials, but also involving a third representation based on so-called Newton polynomials.
 In this terminology, the matrix $U$ performs a basis change from the monomial basis to the Newton basis, and the matrix $L$ performs a basic change from the Newton basis to the evaluation basis.
@@
 The evaluations $P(Z_{\ell})$ can be computed for all $\ell\in\binom{[n]}{\leq D}_{d}$ in time $O^{*}\big(\binom{n}{\leq D}_{d}\big)$.
-For the proof of Theorem 3.1, consider the algorithm TrimmedEvaluation. Throughout we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
+For the proof of Theorem˜3.1, consider the algorithm TrimmedEvaluation. Throughout we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
 Lemma 3.2 (Correctness of TrimmedEvaluation).
@@
 Thus, we correctly return the constant $P$.
-For $n\geq 1$, we write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{0},\dots,P_{d}$ of degrees $D,\dots,D-d$, respectively.
+For $n\geq 1$, we write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many (n1)(n-1)-variate polynomials $P_{0},\dots,P_{d}$ of degrees $D,\dots,D-d$, respectively.
 Formally, each polynomial $P_{i}$ is the coefficient of $P$ viewed as a univariate polynomial in $X_{n}$.
@@
 It is well-known that $V$ is invertible if and only if all grid points $z_{n,i}$ are distinct, and in this case the LU decomposition $V=L\cdot U$ in Line 1 always exists (see e.g.
-[19]).
+[Oruc2000]).
 Further, let $Q_{0},\dotsc,Q_{d}$ as computed in Line 8 of TrimmedEvaluation.
-Note that each polynomial $Q_{j}(X_{1},\dotsc,X_{n-1})=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}(X_{1},\dotsc,X_{n-1})$ has degree $D-j$.
-Indeed, the upper triangular structure of $U$ guarantees that $Q_{j}$ has at most degree $D-j$, because the first non-zero entry in the $j$-th row of $U$ corresponds to the polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
-Calling TrimmedEvaluation$(Q_{j})$, we recursively compute the evaluations $Q_{j}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ in Lines 1 and 1.
-Next, we focus on any iteration $\ell^{\prime}$ of the loop in Line 1.
-Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$, then we show that we correctly compute the evaluations $P(Z_{\ell})$.
-For each $j\in\{0,\dotsc,k\}$, we have
+Note that each polynomial $Q_{j}(X_{1},\dotsc,X_{n-1})=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}(X_{1},\dotsc,X_{n-1})$ has degree DjD-j.
+Indeed, the upper triangular structure of $U$ guarantees that $Q_{j}$ has at most degree DjD-j, because the first non-zero entry in the $j$-th row of $U$ corresponds to the polynomial $P_{j}$ of degree DjD-j while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
+[…]
 $\sum_{i=0}^{k}L_{j,i}\cdot Q_{i}(Z_{\ell^{\prime}})$ $=\sum_{i=0}^{d}L_{j,i}\cdot Q_{i}(Z_{\ell^{\prime}})$
@@
 $=\sum_{i=0}^{d}z_{n,j}^{i}\cdot P_{i}(Z_{\ell^{\prime}})$
-$=P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen,$
-where the first equality follows from the fact that $L_{j,i}=0$ whenever $i>k\geq j$.
-Notably, restricting the computation to $k$ is crucial to ensure that we only use the evaluations $Q_{j}(Z_{\ell^{\prime}})$ that we actually computed.
-This proves that TrimmedEvaluation correctly computes the evaluations $P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen,\dotsc,P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen$ ) in Line 13.
-Since each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ as above, the loop in Line 1 recovers all evaluations $P(Z_{\ell})$.
-As a result, the algorithm TrimmedEvaluation correctly computes all evaluations $P(Z_{\ell})$.
+[…]
+[…]
 Lastly, we prove the running time of TrimmedEvaluation.
@@
 The time to compute the LU decomposition in Line $4$ and the matrix-vector products in Lines $5$ and $9$ can be bounded by $\operatorname{poly}(n,d)$.
-Consequently, the running time of the algorithm without the recursive calls in Line $7$ is $\binom{n}{\leq D}_{d}\cdot M(n,d)$ for some function [^6] In our context, we are content with bounding $M(n,d)=\operatorname{poly}(n,d)$; however, let us remark that following van der Hoeven and Schost [25] we could achieve a dependence on $d$ that is only polylogarithmic.
+Consequently, the running time of the algorithm without the recursive calls in Line $7$ is $\binom{n}{\leq D}_{d}\cdot M(n,d)$ for some function [^6] In our context, we are content with bounding $M(n,d)=\operatorname{poly}(n,d)$; however, let us remark that following van der Hoeven and Schost [HoevenS13] we could achieve a dependence on $d$ that is only polylogarithmic.
 The main insight is that the L- and U-factors of a Vandermonde matrix are sufficiently structured to support matrix-vector operations in time $O(d(\log d)^{O(1)})$.
@@
 Therefore, TrimmedEvaluation runs in time $O^{*}\big(\binom{n}{\leq D}_{d}\big)$.
-Combining Lemma 3.2 and Lemma 3.4 concludes the proof of Theorem 3.1.
+Combining Lemma˜3.2 and Lemma˜3.4 concludes the proof of Theorem˜3.1.
 1
@@
 4if $n=0$ then return the constant $P$.
-5 Compute the LU decomposition $V(z_{n,0},\dots,z_{n,d})^{-1}=L\cdot U$.
-6 for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ do
+5 Compute the UL decomposition $V(z_{n,0},\dots,z_{n,d})^{-1}=U\cdot L$.
+6 for $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$ do
 7   Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
-$\hskip 35.56593pt\begin{pmatrix}\beta_{(\ell^{\prime},0)}\\ \vdots\\ \beta_{(\ell^{\prime},k)}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,k}\\ &\ddots&\vdots\\ &&U_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}\alpha_{(\ell^{\prime},0)}\\ \vdots\\ \alpha_{(\ell^{\prime},k)}\\ \end{pmatrix}.\;$
+$\hskip 35.56593pt\begin{pmatrix}\beta_{(\ell^{\prime},0)}\\ \vdots\\ \beta_{(\ell^{\prime},k)}\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{k,0}&\cdots&U_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}\alpha_{(\ell^{\prime},0)}\\ \vdots\\ \alpha_{(\ell^{\prime},k)}\\ \end{pmatrix}$.\;
 8
 9for $j=0,\dots,d$ do
-10       Recursively call TrimmedInterpolation($\left\lparen(\beta_{(\ell^{\prime},j)})_{\ell^{\prime}}\right\rparen$ ) for ${\ell^{\prime}}\in\binom{[n-1]}{\leq D-j}_{d}$ to interpolate the $(n-1)$-variate polynomial $Q_{j}$.
+[…]
 11
@@
 13Compute the vector of polynomials
-$\hskip 65.44142pt\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{d,0}&\cdots&L_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}.\;$
+$\hskip 65.44142pt\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,d}\\ &\ddots&\vdots\\ &&U_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}$.\;
 14Compute $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$.
-Algorithm 2 TrimmedInterpolation$((\alpha_{\ell})_{\ell})$
+[…]
 3.3 Trimmed Interpolation
-Throughout this section, we show that a polynomial can be (uniquely) interpolated from its evaluations on the trimmed grid points—using essentially the same algorithmic approach as in Section 3.2.
+Throughout this section, we show that a polynomial can be (uniquely) interpolated from its evaluations on the trimmed grid points—using essentially the same algorithmic approach as in Section˜3.2.
 Theorem 3.6 (Trimmed Interpolation).
@@
 The unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ that satisfies $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$ can be computed in time $O^{*}\big(\binom{n}{\leq D}_{d}\big)$.
-Consider the algorithm TrimmedInterpolation for the proof of Theorem 3.6. As before, we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
+Consider the algorithm TrimmedInterpolation for the proof of Theorem˜3.6. As before, we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
 Lemma 3.7 (Correctness of TrimmedInterpolation).
@@
 For the remainder of the proof we assume $n\geq 1$.
+The induction hypothesis implies that we have $Q_{j}(Z_{\ell}^{\prime})=\beta_{(\ell^{\prime},j)}$ as defined in Line 2, for all $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
 Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$.
 It is well-known that $V$ is invertible if and only if the grid points $z_{n,0},\dots,z_{n,d}$ are distinct.
-Additionally, it implies that the LU decomposition in Line 2, namely $V^{-1}=L\cdot U$, always exists (see e.g.
-[19]).
-In the following, we show that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$.
+Moreover, in this case the UL decomposition $V^{-1}=U\cdot L$ from Line 2 always exists (see e.g.
+[Oruc2000]).
+As the matrices $U$ and $L$ are invertible, too, in the following calculation we will use the equivalent formulation $V\cdot U=L^{-1}$.
+We now show that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$.
 Note that each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ with $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ and $\sum_{i=1}^{n-1}\ell^{\prime}_{i}+j\leq D$.
-Fix such a pair $\ell^{\prime}$ and $j$, and let $k=\min\{d,D-\sum_{i=1}^{n-1}\ell^{\prime}_{i}\}$.
-It holds that
-$P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen$ ) $=\sum_{i=0}^{d}z_{n,j}^{i}\cdot P_{i}(Z_{\ell^{\prime}})$
-$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
-$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\beta_{(\ell^{\prime},m)}$
-$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\sum_{h=0}^{k}U_{m,h}\cdot\alpha_{(\ell^{\prime},h)}$
-$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{h=0}^{k}(V^{-1})_{i,h}\cdot\alpha_{(\ell^{\prime},h)}$
-$=\sum_{h=0}^{k}I_{j,h}\cdot\alpha_{(\ell^{\prime},h)}$
-$=\alpha_{(\ell^{\prime},j)},$
-where $I\in\mathbb{F}^{{d+1}\times{d+1}}$ is the identity matrix.
-The last equality uses the fact that $j\leq k$.
+Fix such a pair $\ell^{\prime}$ and $j$.
+As in the algorithm, let $k=\min\{d,D-\sum_{i=0}^{n-1}\ell_{i}^{\prime}\}$.
+Clearly we have $j\leq k$.
+It holds that:
+[…]
+$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}U_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
+$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=i}^{d}U_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
+$=\sum_{m=0}^{d}Q_{m}(Z_{\ell^{\prime}})\cdot\sum_{i=0}^{m}V_{j,i}\cdot U_{i,m}$
+$=\sum_{m=0}^{d}Q_{m}(Z_{\ell^{\prime}})\cdot(L^{-1})_{j,m}$
+$=\sum_{m=0}^{j}Q_{m}(Z_{\ell^{\prime}})\cdot(L^{-1})_{j,m}$
+$=\sum_{m=0}^{j}\beta_{(\ell^{\prime},m)}\cdot(L^{-1})_{j,m}$
+$=\sum_{m=0}^{k}\beta_{(\ell^{\prime},m)}\cdot(L^{-1})_{j,m}$
+$=\alpha_{(\ell^{\prime},j)}.$
 Consequently, TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$.
-As the recursion scheme of TrimmedInterpolation and TrimmedEvaluation are exactly the same, we refer to Lemma 3.4 for a proof of the running time of TrimmedInterpolation.
-This concludes the proof of Theorem 3.6.
+As the recursion scheme of TrimmedInterpolation and TrimmedEvaluation are exactly the same, we refer to Lemma˜3.4 for a proof of the running time of TrimmedInterpolation.
+This concludes the proof of Theorem˜3.6.